# Non-IID data in Federated Learning: A Systematic Review with Taxonomy, Metrics, Methods, Frameworks and Future Directions

### v1

@@ IV-B Interdependence of skew and heterogeneity types
On the other hand, other types of heterogeneity may influence, induce, or exacerbate the problems generated by these core skew types. As an example, running FL training with clients physically located in different places (spatial heterogeneity) can cause data skew between clients, which can further combine with participation skew (e.g., when only recharging mobile devices can take part in training and charging is more common during clients’ night time) to induce notable biases in the global model. Another related concept is temporal heterogeneity, which can make learning harder also in the FL setting even without any data skew.
@@ V Methods and metrics for data heterogeneity in FL
V-A Standardized protocols for dataset partitioning
The term ”partition protocol” describes the systematic processes employed to split a centralized dataset into multiple smaller subsets, thereby creating a federated data landscape. Such protocols for dataset partitioning in FL are crucial for advancing research and practical applications. The current lack of consensus on the sufficient conditions and scenarios for testing solutions to non-IID data has led to a proliferation of partitioning methods and experimental setups. This inconsistency complicates the replication of results, validation of findings, and comparison of different approaches. Establishing agreed-upon methods for creating, quantifying, and evaluating non-IID partitions is essential for enabling fair comparisons, enhancing reproducibility, and providing a common framework for describing various non-IID scenarios.
Leveraged on the exhaustive analysis of the selected papers from the SR, we created a taxonomy for the partition protocols included in the recent literature. Fig. 9 depicts the classification encountered for the methods used in the non-IID research field.
@@ V-A1 Label skew
The Dirichlet partition method is the most employed in research, with 27% of the papers implementing it [114, 115, 116, 117, 118, 119, 72, 120, 121]. It uses the Dirichlet distribution (DD) [122] to allocate samples across clients, simulating non-IID scenarios. A concentration parameter $\alpha$ controls the degree of data heterogeneity. Lower values create more skewed distributions, while higher values approach IID conditions. This method permits generating a range of realistic non-IID datasets for evaluating FL algorithms under varying degrees of data imbalance. The DD generalizes the Beta distribution, which generalizes the Uniform distribution. As a result, it leads to a skewed data split [123].
Another method in this category is the  Generator of Non-IID datasets [124]. It receives as inputs the minimum and maximum number of data points that can be assigned to any partition ($D_{\text{min}}$, $D_{\text{max}}$), the number of labels (or classes) to be sampled for each partition ($L_{\text{num}}$) and the number of partitions ($P$). Then, it creates a set of data partitions $\text{Datas}=(\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{P})$. For each partition $i$, it randomly samples $L_{\text{num}}$ labels from the set of unique categories. It then selects a random number $D_{\text{num}}$ within the range $[D_{\text{min}},D_{\text{max}}]$. A set of weights is sampled from the interval $(0,1)$, normalized by dividing each by the sum of all weights. Each class’s total number of data points is calculated by multiplying $D_{\text{num}}$ by these normalized weights. Finally, data is sampled according to this computed number of data points and the chosen classes, generating the non-IID datasets for each partition.
@@
This method assigns entire classes or subsets of classes to different clients. It has two variations:
@@
One alternative method falling in this category is the  local long-tail (LLT) partition, which simulates Non-IID datasets with a long-tail distribution  [134, 135]. This method divides a global dataset $\mathcal{D}$ with $C$ classes among $K$ clients (where $K=C$). The parameter $\alpha_{l}$ controls the degree of non-IID-ness. For each class $c$, the samples get partitioned into $K$ parts, with one part containing $\alpha_{l}N_{\ast,c}$ samples assigned to a specific client $k$ and the remaining $K-1$ parts containing $\frac{1-\alpha_{l}}{K-1}N_{\ast,y}$ samples distributed to other clients. The latter ensures each client has one dominant class and $C-1$ tail classes. The LLT partition reflects real-world scenarios, like smartphone users favoring certain photo styles, but may not fully capture the distribution characteristics seen in real-world FL tasks. Despite this, the LLT method mirrors imbalanced learning scenarios, serving as a basis for testing data resampling effects in FL.
@@
The most used partition protocol in this category is the Sharding method [136, 137, 138, 139, 140, 141], employed by 20% of the papers. It is a technique used to create non-IID datasets by sorting them based on class labels and then dividing them into smaller subsets, or shards, of equal size. These shards are distributed among clients, each receiving a fixed number. Since each shard contains data from only a few classes, clients have data representing only a small subset of the overall class distribution. This method results in a highly skewed, non-IID distribution, as clients predominantly possess examples from just a few classes, and these do not overlap among the clients. Notice that the case when each client holds only one class is known as Pathological partition [142, 137, 143].
@@ V-A2 Attribute skew
The  clustering method [148] groups the dataset into distinct clusters based on inherent data similarities. These clusters are formed following a non-IID distribution, meaning that the data within each cluster may differ significantly in terms of its features. Once the clustering process is complete, the dataset is divided into subsets corresponding to the clusters, which is helpful for classification or regression tasks. This method ensures that each subgroup reflects a unique data distribution.
@@ V-A3 Quantity skew
The  Min-Size-Dirichlet method [125] helps partition data based on the client’s data size. It corrects the limitation of the Dirichlet for quantity skew method. The algorithm starts by setting an $\alpha$ value for the DD to generate the participation proportions for each client. A minimum required number of examples is then established for each client, with the minimum proportion size defined as $MinSize=\frac{MinRequiredSize}{n}$, where $n$ represents the total number of examples in the centralized dataset. If any calculated proportions are smaller than $MinSize$, they are replaced by $MinSize$. Finally, the proportions are normalized between 0 and 1. The latter ensures the method will converge even when the dataset is small, or the number of clients is significant.
@@ V-B Non-IIDness quantification metrics
The term ”non-IID metric” is used to quantify how much a dataset deviates from being IID, measuring the degree of data heterogeneity. Although research on quantifying non-IIDness in FL is limited, its importance cannot be overstated for several relevant reasons. First, it provides a clear understanding of the statistical diversity among participating clients, directly affecting model convergence, stability, and overall performance [154]. Accurate measurements for measuring non-IIDness help researchers and practitioners anticipate challenges in training, allowing for the design of tailored algorithms to address the specific level of heterogeneity present. These metrics also enable standardized evaluations and benchmarking of FL methods across different datasets and scenarios [155]. Furthermore, they guide the development of standardized testbeds that resemble real-world FL scenarios more accurately to produce robust and generalizable FL solutions.
Based on the comprehensive review of papers selected in our SR, we developed a classification for non-IID metrics in recent literature. Fig. 10 illustrates our taxonomy of measures used to quantify non-IID characteristics in FL research.
@@ V-B1 Label skew
These metrics compare label distributions between clients or against a global distribution. Label distributions refer to the proportion or frequency of different classes within a dataset. The comparison uses widely known distances and divergences like Earth Mover’s distance (also known as Wasserstein distance) [156, 157, 140, 158], Hellinger distance [125, 159], Jensen-Shannon divergence [160, 161], Kullback-Leibler divergence [162], and similarities like the Cosine [37, 143] and Jaccard [163].
@@ V-B2 Attribute skew
These metrics measure the differences in feature distributions across clients. This category includes only metrics using the attributes distribution per client.
Distribution-based
These metrics assess attribute distribution differences using measures like Hellinger distance, Earth mover’s distance, and Jensen-Shannon divergence, applied to feature spaces rather than label spaces [125].
@@ V-B3 Quantity skew
The metrics evaluate the imbalance in data volume across clients. This category includes only metrics leveraged on the client’s participation distribution.
Distribution-based
These metrics analyze the distribution of data quantities across clients, often using similar measures as in label and attribute skew (e.g., Hellinger distance, Earth mover’s distance, Jensen-Shannon divergence) but applied to sample counts or proportions [125].
@@ V-B4 Label + Attribute skew
The other two metrics proposed in the literature are inner variation and outer variation [152]. inner variation measures the randomness in sampling data from a client’s local distribution $P_{i}$ and is defined as $\mathbb{E}[g_{i}]-\mathbb{E}[g_{i}^{2}]$, where $g_{i}$ is the gradient of client $i$. Outer variation captures the difference between a client’s local distribution $P_{i}$ and the global distribution $\frac{1}{|H|}\sum_{i\in H}P_{i}$, and is given by $\mathbb{E}[g_{i}]-\mathbb{E}[u^{2}]$, where $u$ represents the global gradient. In IID settings, outer variation is zero, while in feature and label-skewed scenarios, it becomes non-zero, indicating higher non-IIDness.
@@
where $\mathcal{D}=\bigcup_{i=1}^{K}\mathcal{D}_{i}$, $\mathcal{D}_{i}^{k}$ denotes the data belonging to the $k$-th class in $\mathcal{D}_{i}$. $|C_{i}|$ is the number of classes in $\mathcal{D}_{i}$. ()($\cdot$)( ⋅ ) represents the first order moment, $\sigma(\cdot)$ is the standard deviation and is used to normalize the scale, and $\|\cdot\|_{2}$ indicates the $\ell_{2}$-norm. The intuition behind Eq. 10 is to measure the distance between the average data representations from different classes in feature space on a given client and the counterpart over all the other clients.
@@ VI Non-IID solutions in FL
This section briefly discusses the leading solutions for non-IID datasets in FL to tackle the downstream effects exposed in Section II. Notice that most of these solutions focus on one type of non-IIDness, which is something future research should improve. In particular, we describe in the following paragraphs the solutions that are employed the most in the selected papers of our SR (see Fig. 11).
@@
Dynamic Client Selection [37, 192, 157, 39, 193, 194]. Most FL papers assume unbiased client participation, where clients are selected randomly or in proportion to their data sizes. In Dynamic Client Selection, clients are selected employing a bias strategy. In other words, clients are dynamically chosen at each round (or epoch) to achieve a good trade-off between convergence speed and solution bias. Its advantages include faster convergence by prioritizing clients with higher-quality data, better handling system heterogeneity, and efficient use of resources. However, it may introduce selection bias, leading to over-reliance on specific clients, reduced fairness, and decreased model generalization. Additionally, it can increase the complexity of client selection, add communication overhead, and risk underrepresented clients with limited computational resources.
@@ VII Frameworks and tools for heterogeneous data in FL
VIII Lessons learned
This section highlights the essential findings and critical insights obtained through our SR of the current literature on non-IID data in FL. We aim to provide practitioners with robust conclusions and thought-provoking considerations for developing practical solutions in this challenging area of FL. These lessons encompass various aspects, including the quantification and classification of non-IID-ness, the interplay between different types of data skew, and the practical implications for implementation and standardized framework selection.
@@ VIII Lessons learned
    Prevalence of label skew studies. Our SR indicates that a majority of papers focus on label skew. While this is an important aspect of non-IID data, causing the major deterioration of FL models [222, 223, 224, 225, 226, 227], the disproportionate attention may lead to overlooking other critical types of non-IID skews.
@@
    Few works in multimodality skew research. The multimodality skew, as exposed in Section IV, is a topic that has not been formally included in the previous surveys or systematic reviews for non-IIDness. Despite its relevance, the research papers that include that topic are just a few [213, 86, 228, 229, 84, 230, 85, 231]. As real-world FL applications often involve diverse data types and modalities, understanding and addressing multimodality skew is a key aspect for practical implementations with the aim of understanding and tackling the consequences of non-IID data in FL.
@@ IX Future directions and trends
    Multimodal learning. A promising avenue for future research in FL is the exploration of multimodal learning techniques to address the increasingly complex nature of real-world data. As applications of FL expand across diverse domains, there is a growing need to develop robust methods capable of handling multiple data modalities simultaneously, such as text, images, audio, and sensor data [241, 87]. Future research should focus on creating novel FL algorithms that can effectively integrate and learn from these heterogeneous data types while preserving privacy and maintaining efficiency. This direction presents unique challenges in the context of non-IID data, as different modalities may exhibit varying degrees and types of heterogeneity across clients.
@@
    Addressing data heterogeneity impact. Addressing data heterogeneity remains a central challenge in FL, necessitating innovative approaches to enhance model performance and adaptability in non-IID settings [242, 76, 243]. Future research should prioritize the development of more effective algorithms designed to handle diverse forms of non-IIDness, specifically under high levels of non-IIDness. These advanced algorithms should aim to mitigate the negative impacts of non-IID data on model convergence, fairness, and generalization (robustness).

### v2

@@ IV-B Interdependence of skew and heterogeneity types
On the other hand, other types of heterogeneity may influence, induce, or worsen the problems generated by these core skew types. As an example, running FL training with clients physically located in different places (spatial heterogeneity) can cause data skew between clients, which can further combine with participation skew (e.g., when only recharging mobile devices can take part in training and charging is more common during clients’ night time) to induce notable biases in the global model. Another related concept is temporal heterogeneity, which can make learning harder also in the FL setting even without any data skew.
@@ V Methods and metrics for data heterogeneity in FL
V-A Standardized dataset’s partition protocols
Protocols for dataset partitioning in FL are crucial for advancing research and practical applications. The current lack of consensus on the sufficient conditions and scenarios for testing solutions to non-IID data has led to an increase of partitioning protocols and experimental setups. This inconsistency complicates the replication of results, validation of findings, and comparison of different approaches. Establishing agreed-upon methods for creating, quantifying, and evaluating non-IID partitions is essential for enabling fair comparisons, enhancing reproducibility, and providing a common framework for describing various non-IID scenarios.
Leveraged on the exhaustive analysis of the selected papers from the survey, we created a taxonomy for the partition protocols included in the recent literature. Fig. 9 depicts the classification encountered for the methods used in the non-IID research field.
@@ V-A1 Label skew
The Dirichlet partition protocol is the most employed in research, with 27% of the papers implementing it [117, 118, 119, 120, 121, 122, 75, 123, 124]. It uses the Dirichlet distribution (DD) [125] to allocate samples across clients, simulating non-IID scenarios. A concentration parameter $\alpha$ controls the degree of data heterogeneity. Lower values create more skewed distributions, while higher values approach IID conditions. This method permits generating a range of realistic non-IID datasets for evaluating FL algorithms under varying degrees of data imbalance. The DD generalizes the Beta distribution, which generalizes the Uniform distribution. As a result, it leads to a skewed data split [126].
Another method in this category is the  Generator of Non-IID datasets [127]. It receives as inputs the minimum and maximum number of data points that can be assigned to any partition ($D_{\text{min}}$, $D_{\text{max}}$), the number of labels (or classes) to be sampled for each partition ($L_{\text{num}}$) and the number of partitions ($P$). Then, it creates a set of data partitions $\text{Dataset}=(\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{P})$. For each partition $i$, it randomly samples $L_{\text{num}}$ labels from the set of unique categories. It then selects a random number $D_{\text{num}}$ within the range $[D_{\text{min}},D_{\text{max}}]$. A set of weights is sampled from the interval $(0,1)$, normalized by dividing each by the sum of all weights. Each class’s total number of data points is calculated by multiplying $D_{\text{num}}$ by these normalized weights. Finally, data is sampled according to this computed number of data points and the chosen classes, generating the non-IID datasets for each partition.
@@
This method assigns entire classes or datasets of classes to different clients. It has two variations:
@@
One alternative method falling in this category is the  local long-tail (LLT) partition, which simulates Non-IID datasets with a long-tail distribution  [137, 138]. This method divides a global dataset $\mathcal{D}$ with $C$ classes among $K$ clients (where $K=C$). The parameter $\alpha_{l}$ controls the degree of non-IID-ness. For each class $c\in C$, the samples get partitioned into $K$ parts, with one part containing $\alpha_{l}N_{\ast,c}$ samples assigned to a specific client $k$ and the remaining $K-1$ parts containing $\frac{1-\alpha_{l}}{K-1}N_{\ast,y}$ samples distributed to other clients. The latter ensures each client has one dominant class and $C-1$ tail classes. The LLT partition reflects real-world scenarios, like smartphone users favoring certain photo styles, but may not fully capture the distribution characteristics seen in real-world FL tasks. Despite this, the LLT method mirrors imbalanced learning scenarios, serving as a basis for testing data resampling effects in FL.
@@
The most used partition protocol in this category is the Sharding method [139, 140, 141, 142, 143, 144], employed by 20% of the papers. It is a technique used to create non-IID datasets by sorting them based on class labels and then dividing them into smaller datasets (a.k.a shards) of equal size. These shards are distributed among clients, each receiving a fixed number. Since each shard contains data from only a few classes, clients have data representing only a small subset of the overall class distribution. This method results in a highly skewed, non-IID distribution, as clients predominantly possess samples from just a few classes, and these do not overlap among the clients. Notice that the case when each client holds only one class is known as Pathological partition [145, 140, 146].
@@ V-A2 Attribute skew
The  clustering method [151] groups the dataset into distinct clusters based on inherent data similarities. These clusters are formed following a non-IID distribution, meaning that the data within each cluster may differ significantly in terms of its features. Once the clustering process is complete, the dataset is divided into datasets corresponding to the clusters, which is helpful for classification or regression tasks. This method ensures that each subgroup reflects a unique data distribution.
@@ V-A3 Quantity skew
The  Min-Size-Dirichlet method [128] helps partition data based on the client’s data size. It corrects the limitation of the Dirichlet for quantity skew method. The algorithm starts by setting an $\alpha$ value for the DD to generate the participation proportions for each client. A minimum required number of samples is then established for each client, with the minimum proportion size defined as $MinSize=\frac{MinRequiredSize}{n}$, where $n$ represents the total number of samples in the centralized dataset. If any calculated proportions are smaller than $MinSize$, they are replaced by $MinSize$. Finally, the proportions are normalized between 0 and 1. The latter ensures the method will converge even when the dataset is small, or the number of clients is significant.
@@
From Fig. 9, we notice that partition protocols simulating and controlling two or three types of data skew simultaneously do not appear in the papers reviewed. This observation reveals a significant gap in current partitioning approaches, as real-world FL scenarios present combinations of different skew types concurrently.
@@ V-B Non-IIDness quantification metrics
The term ”non-IID metric” is used to quantify how much a dataset deviates from being IID, measuring the degree of data heterogeneity. Although research on quantifying non-IIDness in FL is limited, its importance cannot be overstated for several relevant reasons. First, it provides a clear understanding of the statistical diversity among participating clients, directly affecting model convergence, stability, and overall performance [157]. Accurate measurements for measuring non-IIDness help researchers and practitioners anticipate challenges in training, allowing for the design of tailored algorithms to address the specific level of heterogeneity present. These metrics also enable standardized evaluations and benchmarking of the robustness of FL methods across different datasets and non-IID scenarios [158]. Furthermore, they guide the development of standardized testbeds that resemble real-world FL scenarios more accurately to produce robust and generalizable FL solutions.
Based on the comprehensive review of papers selected in our survey, we developed a classification for non-IID metrics in recent literature. Fig. 10 illustrates our taxonomy of measures used to quantify non-IID characteristics in FL research.
@@ V-B1 Label skew
These metrics compare label distributions between clients or against a global distribution. Label distributions refer to the proportion of different classes within a dataset. The comparison uses widely known distances and divergences like Earth Mover’s distance (also known as Wasserstein distance) [159, 160, 143, 161], Hellinger distance [128, 162], Jensen-Shannon divergence [163, 164], Kullback-Leibler divergence [165], and similarities like the Cosine [40, 146] and Jaccard [166].
@@ V-B2 Attribute skew
These metrics measure the differences in feature distributions across clients. This category includes only metrics using the attributes distribution per client (Distribution-based). It includes measures like Hellinger distance, Earth mover’s distance, and Jensen-Shannon divergence, applied to feature spaces rather than label spaces [128].
@@ V-B3 Quantity skew
The metrics evaluate the imbalance in data volume across clients. This category includes only metrics leveraged on the client’s participation distribution (Distribution-based). It often uses measures similar to label and attribute skew (e.g., Hellinger distance, Earth mover’s distance, Jensen-Shannon divergence) but applied to sample counts or proportions [128].
@@ V-B4 Label + Attribute skew
The other two metrics proposed in the literature are inner variation and outer variation [155]. Inner variation measures the randomness in sampling data from a client’s local distribution $f_{Y}^{(i)}$ and is defined as $\mathbb{E}[g_{i}]-\mathbb{E}[g_{i}^{2}]$, where $g_{i}$ is the gradient of client $i$. Outer variation captures the difference between a client’s local distribution $P_{i}$ and the global distribution $\frac{1}{|H|}\sum_{i\in H}P_{i}$, and is given by $\mathbb{E}[g_{i}]-\mathbb{E}[u^{2}]$, where $u$ represents the global gradient. In IID settings, outer variation is zero, while in feature and label-skewed scenarios, it becomes non-zero, indicating higher non-IIDness.
@@
where $\mathcal{D}=\bigcup_{i=1}^{K}\mathcal{D}_{i}$, $\mathcal{D}_{i}^{k}$ denotes the data belonging to the $k$-th class in $\mathcal{D}_{i}$. $|C_{i}|$ is the number of classes in $\mathcal{D}_{i}$, $\sigma(\cdot)$ is the standard deviation and is used to normalize the scale, and $\|\cdot\|_{2}$ indicates the $\ell_{2}$-norm. The intuition behind Eq. 10 is to measure the distance between the average data representations from different classes in feature space on a given client and the counterpart over all the other clients.
@@ V-B5 Label + Attribute + Quantity skew
Our survey reveals a notable absence of metrics that can simultaneously measure combinations of data heterogeneities, such as label with quantity skew, attribute with quantity skew, or combinations with spatiotemporal heterogeneity, participation skew, and modality skew. This measurement gap is particularly concerning, given that these heterogeneities co-exist in real-world FL scenarios. Without combined metrics, researchers and practitioners lack the tools to effectively quantify and understand how multiple types of data heterogeneity jointly impact model performance.
@@ VI Non-IID solutions in FL
This section briefly discusses the leading solutions for non-IID datasets in FL to tackle the downstream effects exposed in Section II. Notice that most of these solutions focus on one type of non-IIDness, which is something future research should improve. In particular, we describe in the following paragraphs the solutions that are employed the most in the selected papers of our survey (see Fig. 11).
@@
Dynamic Client Selection [40, 195, 160, 42, 196, 197]. Most FL papers assume unbiased client participation, where clients are selected randomly or in proportion to their data sizes. In Dynamic Client Selection, clients are selected employing a bias strategy. In other words, clients are dynamically chosen at each round to achieve a good trade-off between convergence speed and solution bias. Its advantages include faster convergence by prioritizing clients with higher-quality data, better handling system heterogeneity, and efficient use of resources. However, it may introduce selection bias, leading to over-reliance on specific clients, increased bias, and decreased model generalization. Additionally, it can increase the complexity of client selection, add communication overhead, and risk underrepresented clients with limited computational resources.
@@ VII Frameworks and tools for heterogeneous data in FL
VIII Essential findings and key insights
This section highlights the essential findings and critical insights obtained through our survey of the current literature on non-IID data in FL. We aim to provide practitioners with robust conclusions and thought-provoking considerations for developing practical solutions in this challenging area of FL. These lessons encompass various aspects, including the quantification and classification of non-IID-ness, the interplay between different types of data skew, and the practical implications for implementation and standardized framework selection.
@@ VIII Essential findings and key insights
    Prevalence of label skew studies. Our survey indicates that a majority of papers focus on label skew. While this is an important aspect of non-IID data, causing the major deterioration of FL models [225, 226, 227, 228, 229, 230], the disproportionate attention may lead to overlooking other critical types of non-IID skews.
@@
    Few works in multimodality skew research. The multimodality skew, as exposed in Section IV, is a topic that has not been formally included in the previous surveys for non-IIDness. Despite its relevance, the research papers that include that topic are just a few [216, 89, 231, 232, 87, 233, 88, 234]. As real-world FL applications often involve diverse data types and modalities, understanding and addressing multimodality skew is a key aspect for practical implementations with the aim of understanding and tackling the consequences of non-IID data in FL.
@@ IX Future directions and trends
    Modality skew research. A promising avenue for future research in FL is the exploration of multimodal learning techniques to address the increasingly complex nature of real-world data. As applications of FL expand across diverse domains, there is a growing need to develop robust methods capable of handling multiple data modalities simultaneously, such as text, images, audio, and sensor data [244, 90]. Future research should focus on creating novel FL algorithms that can effectively integrate and learn from these heterogeneous data types while preserving privacy and maintaining efficiency. This direction presents unique challenges in the context of non-IID data, as different modalities may exhibit varying degrees and types of heterogeneity across clients.
@@
    Addressing data heterogeneity impact. Addressing data heterogeneity remains a central challenge in FL, necessitating innovative approaches to enhance model performance and adaptability in non-IID settings [245, 79, 246]. Future research should prioritize the development of more effective algorithms designed to handle diverse forms of non-IIDness, specifically under high levels of non-IIDness. These advanced algorithms should aim to mitigate the negative impacts of non-IID data on model convergence, bias, and generalization (robustness).
@@
  7. 7.
    Public datasets with non-IIDness quantification. A crucial step toward advancing research on non-IID data in FL is the creation and dissemination of publicly available datasets with explicit and standardized characterizations of their non-IID properties. Including clear non-IID metrics to quantify the extent and types of non-IIDness in these datasets would facilitate benchmarking, reproducibility, and the development of more robust methods tailored to diverse real-world scenarios.